OpenStudy (anonymous):

write the standard form of the equation of the parabola that has a vertex of (6,3) and passes through the point (-4,-8) i think the answer is f(x) = -11/100(x-6)^2+3 plz help

OpenStudy (tkhunny):

1) That would be the Vertex Form, not Standard Form. 2) Is it right? How did you get that?

OpenStudy (anonymous):

out of the answer choice i pick that one i completely guess. im really confused with this question

OpenStudy (tkhunny):

First Problem, there are TWO such parabolas as the problem has been presented. Which one shall we pick?

OpenStudy (anonymous):

(6,3)

OpenStudy (tkhunny):

?? The vertex is (6,3) in both solutions. y-3 = a(x-6)^2 OR x-6 = b(y-3)^2 Both can be solved to pass through (-4,-8).

OpenStudy (anonymous):

im so confused

OpenStudy (tkhunny):

Do those two things I have written look like parabolas?

OpenStudy (anonymous):

yes they look like parabolas

OpenStudy (anonymous):

i give up im just going to guess. thank you for the help i really appreciate it

OpenStudy (tkhunny):

It's a shame. Just substitute the values.